find x:
2 sqrt of x-3 = 12

Answer by nerdybill(7384)

2 sqrt(x-3) = 12
Begin by dividing both sides by 2:
sqrt(x-3) = 6
Now, square both sides:
x-3 = 36
Finally, add 3 to both sides:
x = 39
.
.
To check, plug it back in to the original equation:
2 sqrt(x-3) = 12
2 sqrt(39-3) = 12
2 sqrt(36) = 12
2(6) = 12
12 = 12 (checks)